OpenStudy (anonymous):

there are 12 books on algebra and calculus in a library ,the books on the same subject being different . if the numbr of selections each of whish consists 3 buks on each topic is greatest, then the numbr of books of algebra and calculus in the library is......... a.3 & 9 b.4 & 8 c.5 & 7 d..6 & 6

OpenStudy (yuki):

Since the group of 3 books can be found by combination, why don't we use that ?

OpenStudy (yuki):

for example, if algebra was 3 and calculus was 9, then the total number of ways that you can chose from those would be 3 * 9 = 27.

OpenStudy (yuki):

because there are 3 ways to choose an alg. book and 9 ways to choose a calc. book.

OpenStudy (yuki):

so among a,b,c,d in your choice, you just choose the one that has the biggest value.

OpenStudy (anonymous):

whats the value of log5 512

OpenStudy (anonymous):

without using calculator..

OpenStudy (anonymous):

conditions are give like this as...log 2 = 0.3010 and log 3 = 0.47771

OpenStudy (anonymous):

so whats the value of log base5 512.?

OpenStudy (yuki):

why don't you use the fact that 512 = 2^9 ?

OpenStudy (yuki):

so log_5 (2^9) = 9 log_5 (2)

OpenStudy (yuki):

if you know how to use the change of base then that's what you will do next

OpenStudy (anonymous):

idk to change the base....m the student preparing for medicine i dnt have much knowledge in math.:)

OpenStudy (yuki):

the change of base looks like this \[loga(b) = (logc(b))/(logc(a))\]

OpenStudy (yuki):

c is any number you would like it to be, in your case the base seems to be 10 since log(2) is given. Just to let you know, we also need a log(5) as well, so make sure that your given numbers have it. If not, you can't solve this one.

OpenStudy (anonymous):

and the option is given like this a.2.70 b.2.967 c.3.876 d.3.912

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

so is option no c is correct.